August 20, 2013. The TEM camera computer is repaired and the TEM is back in service.
Currently browsing category
August 16, 2013. The computer that runs the TEM camera is having trouble. I took it to the shop for repair and we are out of commission this afternoon. Hopefully back online later this afternoon or Monday.
July 10, 2013. We will have to begin using liquid nitrogen in the TEM cold trap. No new users has seen the done so I’ll have to show you individually. Just come by my office.
July 9, 2013. I “found” the electron beam and dialed in a good alignment on the TEM today and it is working fine. You can use it now. Note that the lab temp and humidity are such that we will begin using the liquid nitrogen cold traps. I will show …
July 8, 2013. The TEM lab air conditioning failed on Saturday and the lab to heated up over 85 F. As a result the lenses shutdown. The AC is back on now and I have fans running to direct cool air into the labs and onto the lens PS. Bottom …
July 3, 2013. The TEM is still not working. I will work on it again this afternoon. July 1, 2013. The TEM is down. Evidently something happened after the power outage last Wednesday night. I’m working on it now and will contact you again with updates.
June 27, 2013. The TEM is not functioning properly and will be unavailable until Owen returns next week. We are sorry for any inconvenience.
June 6, 2013. The TEM is repaired and, after the gun vacuum improves overnight, should be ready for the first appointment (Yang). Note that one of the standard sample holders was bent somehow. I’ll send it in for repair. Be careful – repairs like this lead to higher use charges.
June 5, 2013. I think I know what is wrong and I will start on the fix tomorrow morning. It definitely won’t be ready for the tomorrow morning appointment (Yifei) but may be operational by afternoon. Look for a status update ~noon.
June 4, 2013. Unfortunately, the TEM did not start and cannot be used. I have called the engineer and hope to hear from him early tomorrow.